OpenStudy (anonymous):

did you guys know that 21^2 -21 = 420. Just found out. lol

OpenStudy (alexandervonhumboldt2):

i know this.

OpenStudy (♪chibiterasu):

Why is it such a big deal.

OpenStudy (alexandervonhumboldt2):

11^2-11=110 what is interesting in it?

OpenStudy (alexandervonhumboldt2):

10^2-10=90

hartnn (hartnn):

21 (21 -1) = 21* 20 = 420 thats how i did it in my head

OpenStudy (anonymous):

420 joke obviously

TheSmartOne (thesmartone):

@hartnn teach me more tricks and tips ;p

OpenStudy (anonymous):

lol.

OpenStudy (chantysquirrel1129**):

hahaha wow cx ._.